a 6 volt battery is connected to a 0.3 m wire with a diameter of 1.0mm has a current of 5. ampere flows through the wire. How much current will flow if the wire is replaced by 1.20 m wire of the same wire and diameter?



1
Expert's answer
2021-06-21T16:26:41-0400

Gives

Battery (V)=6V

Length of wire ("{l_1})=0.3m"

Daimeter (d1)=1 mm

Diameter (d2)=1mm

Current ("I_1" )=5A

Length"(l_2)=1.20m"

Area both wire are same

"R=\frac{\rho l}{A}"

"V=IR"

"V=I\frac{\rho l}{A}"

Similarly we can represent

Both wire relationship


"\frac{I_1l_1}{A_1}=\frac{I_2l_2}{A_2}"

"I_2=\frac{l_1I_1}{A_1} \times\frac{A_2}{l_2}"

Put value


"I_2=\frac{5\times0.3}{A}\times\frac{A}{1.20}=1.25A"